OpenStudy (anonymous):

When he goes out to eat, Austin always leaves a tip for his server that is equal to 15% of his bill. Let f = the amount of the bill. Which algebraic expression could be used to calculate Austin's total cost of eating out at any restaurant?

OpenStudy (helder_edwin):

if \(f\) is the amount of the bill and he leaves a 15% tip then the total cost would be \[ \large C=f+f\cdot15\%=f\cdot(1+0.15)=1.15f \]

OpenStudy (anonymous):

So f.15 would be the algebraic expression?

OpenStudy (helder_edwin):

1.15f